OpenStudy (shamil98):

Solve the polynomial equation. f(x) = x^3 - 4x^2 + x + 6 How do I do this? ..

hartnn (hartnn):

solve means? do you need to find roots/zeros ?

OpenStudy (shamil98):

I think so.

OpenStudy (shamil98):

The instructions are as above..

OpenStudy (shamil98):

A bit vague though.

hartnn (hartnn):

which topic does this belong to ? can we use rational root theorem ? heard about it ?

OpenStudy (shamil98):

Pre-cal, polynomial and rational functions. Yeah, the rational root theorem is allowed.

hartnn (hartnn):

good, so you know about it ? the roots can be factors of \(\pm\)6/1 start by test 1, then -1 is any of these a root of f(x) ? do you know how to test ?

OpenStudy (shamil98):

I have no idea what it is . Testing would just be substituting the value into the function right?

hartnn (hartnn):

yes test +1 first

hartnn (hartnn):

if you get a 0, then it is a root

OpenStudy (shamil98):

f(x) = x^3 - 4x^2 + x + 6 f(1) = 1 - 4(1) + 1 + 6 f(1) = 1-4 + 1 + 6 f(1) = -3 + 7 = 4 so it isn't?

hartnn (hartnn):

correct now test -1

OpenStudy (ranga):

Jackpot in 3.... 2... 1....

hartnn (hartnn):

lol

OpenStudy (shamil98):

lol

OpenStudy (shamil98):

f(-1) = -1^3 - 4(-1)^2 + -1 + 6 f(-1) = -1 - 4 - 1 + 6 = -6 + 6 = 0

OpenStudy (shamil98):

so that's one.

OpenStudy (shamil98):

finding all of these must be a pain.

hartnn (hartnn):

yes now you can continue doing this for +2,-2, +3,-3 .... but if you know synthetic division, it'll simplify things

OpenStudy (shamil98):

I was suppose to learn that.. never got around to it. Care to enlighten? lol.

OpenStudy (ranga):

http://www.youtube.com/watch?v=bZoMz1Cy1T4

OpenStudy (shamil98):

ranga the video helps, but what am I suppose to divide it by? it's not a rational function

hartnn (hartnn):

its not ? how ? since we have -1 as on root, take -1 on left (similar to polynomial division by x+1)

OpenStudy (shamil98):

so you would have to do that for each possible value tho..

OpenStudy (shamil98):

seems a bit tedious lol.

hartnn (hartnn):

each ? do it for -1 and you directly get a simple quadratic equation

hartnn (hartnn):

and you can solve quadratic equations right ?

OpenStudy (shamil98):

yes lol.

hartnn (hartnn):

problem solved :P

OpenStudy (shamil98):

 so i tried synthetic division